For the following 4 options (one must be chosen), what is the smallest MARR for the Do Nothing option to be chosen? Option
Answer #1

At first, we will calculate the IRR of the option A, B and C

Let's say, it is rA, rB and rC

Now, we know NPV of cash ct at time t is ct/(1+r)^t where r is the interest rate, using it, we have

1000=153/(1+rA)+153/(1+rA)^2+...+153/(1+rA)^10 = (153/(1+rA))*(1-1/(1+rA)^10)/(1-1/(1+rA)) =(153/rA)*(1-1/(1+rA)^10)

This gives us rA=8.6%

Now, 1500=(207/rB)*(1-1/(1+rB)^10), this gives us rB=6.4%

Now, 2000=(261/rC)*(1-1/(1+rC)^10), this gives us rC=5.26%

So, the smallest of them is 5.26% which will be the smallest MARR

Thus, the correct option is A